OpenStudy (rebeccapink6398):

@31356 plz help! :D So I have another question.. Kyle paid $96 in interest to borrow $7200 for 4 months. What annual interest rate did Kyle pay? A. 1.33% B. 4% C. 5.33% D. 16% so would I take the $7200*..? and that's where im stuck

OpenStudy (31356):

Let me see........

OpenStudy (31356):

Hmm..........

OpenStudy (rebeccapink6398):

I gave u a medal now..so I would not forget lol :D

OpenStudy (31356):

Thanks! lol :D

OpenStudy (rebeccapink6398):

lol yeah :D

OpenStudy (31356):

So you have to change $96 into a percentage first.

OpenStudy (rebeccapink6398):

ok so that is.. 0.096??

OpenStudy (rebeccapink6398):

or.. 9600% ?

OpenStudy (31356):

Oh sorry, I was helping my brothers.

OpenStudy (rebeccapink6398):

lol..GOOD SISTER! lol

OpenStudy (31356):

Thanks :D

OpenStudy (31356):

Here, she can help, since I have to help my brothers for a while. :D @MelissaHolmes

OpenStudy (rebeccapink6398):

ok :D

OpenStudy (anonymous):

so 96 is just a number...

OpenStudy (31356):

Yea, $

OpenStudy (31356):

Or dollars, lol

OpenStudy (rebeccapink6398):

yes

OpenStudy (31356):

Or ask @whpalmer4 for more help?

OpenStudy (whpalmer4):

\[I = Prt\]\[P = 7200\]\[I = 96\]\[t=4/12\]\[I=Prt\]\[\frac{I}{Pt} = r\]\[r = \frac{96}{7200*(4/12)} = \]

OpenStudy (whpalmer4):

The value of \(r\) will be a decimal, which you will convert to a percentage by multiplying with 100%